SCALING IN A MAP OF THE TWO-TORUS

KEITH M. BRIGGS AND GONZALO ÁLVAREZ

Abstract. We discuss scaling in the parameter space of a family of maps arising from the iteration of a map of the two-torus defined in terms of a Jacobian elliptic function. This map appears to show a complex analog of the Feigenbaum–Kadanoff–Shenker scaling found in bifurcation sequences of circle maps.